### v1

# Overcoming Output Dimension Collapse: How Sparsity Enables Zero-shot Brain-to-Image Reconstruction at Small Data Scales

Abstract
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. Achieving such reconstruction with limited training data requires generalization beyond the training set, a task known as zero-shot prediction. Despite its importance, we still lack theoretical guidelines for achieving efficient and accurate reconstruction. In this paper, we provide a theoretical analysis of two widely used models for translating brain activity to latent image features. We define the data scale as the ratio of the number of training samples to the latent feature dimensionality and characterize how each model behaves across data scales. We first show that the naive linear regression model, which uses a shared set of input variables for all outputs, suffers from “output dimension collapse” at small data scales, restricting generalization beyond the training data. We then mathematically characterize the prediction error of the sparse linear regression model by deriving formulas linking prediction error with data scale and other problem parameters. Leveraging the sparsity of the brain-to-feature mapping, this approach enables accurate zero-shot prediction even at small data scales without trapping in output dimension collapse. Our results provide a theoretical guideline for achieving zero-shot reconstruction and highlight the benefits of variable selection in brain-to-image reconstruction.
keywords:
Brain decoding , Brain-to-image reconstruction , Zero-shot prediction , Sparse linear regression
[…]
1 Introduction
Advances in brain-to-image reconstruction are enabling us to externalize the subjective visual experiences encoded in the brain as images. To uncover neural representations and move toward practical applications in medicine and industry, it will be necessary to reconstruct a broad range of subjective visual experiences. Nevertheless, a fundamental constraint governs this task: the space of possible visual stimuli is vast, yet we can only collect a limited number of brain-image pairs for training. To capture the full spectrum of subjective experiences, a model must be able to predict previously unseen stimuli, namely zero-shot prediction (Shirakawa et al., 2025). Despite its importance, we still lack theoretical guidelines for achieving accurate reconstruction under these constraints.
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1). A translator module converts brain activity into latent features that represent the image content, and a generator module transforms these features into reconstructed images. This paper focuses on the translator module, trained on a limited pairs of brain activity and latent features. We quantify this limitation by the data scale, the ratio of the number of training pairs to the dimensionality of the latent features. To realize zero-shot reconstruction, the translator must be able to predict valid latent features for unseen stimuli at small data scales.
Figure 1: Translator–Generator pipeline. The translator converts brain activity into latent features, and the generator transforms latent features into reconstructed images. A naive multivariate linear regression model and a sparse multivariate linear regression model are commonly used as translators. Whereas a naive model uses a single shared set of input variables to predict all output dimensions, a sparse model allows each output dimension to select its own subset of input variables.
Two linear regression models are widely used as translators: naive multivariate linear regression and sparse multivariate linear regression. A naive multivariate linear regression model, exemplified by ridge regression, uses a single shared set of input variables to predict all output dimensions and has been adopted in several high-profile studies (Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). This strategy is straightforward and low-cost, yet Shirakawa et al. (2025) reports that such models fail to generalize beyond the training domain. They attribute this failure to “output dimension collapse,” in which predictions are confined to the span of the training outputs under dataset bias. In contrast, a sparse multivariate linear regression model allows each output dimension to select its own subset of input variables. Earlier work adopted sparse regression as the translator, assuming that “brain-like” latent features induce a sparse brain-to-feature mapping (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019).
In this paper, we provide a theoretical analysis of translator prediction under small data scales. First, complementing prior work that emphasizes dataset bias (Shirakawa et al., 2025), we show that small data scales cause output dimension collapse in naive multivariate linear regression. Using the mathematically derived best prediction under this collapse, we evaluate its impact on reconstruction pipelines that employ naive translators. Second, we examine how sparsity can enable accurate prediction at small data scales by characterizing the prediction error of naive and sparse translators. Our contributions are as follows:
  • 1.
    We show that reconstruction pipelines using naive linear regression suffer from output dimension collapse at small data scales.
  • 2.
    We mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters.
  • 3.
    We demonstrate that sparse linear regression can escape output dimension collapse and achieve accurate prediction at small data scales by leveraging sparse brain-to-feature structure.
2 Preliminaries
Many recent reconstruction methods employ a typical architecture, referred to as the Translator–Generator pipeline (Shirakawa et al., 2025) (Figure 1). In this framework, a module called the translator converts brain activity into latent features, and a module called the generator transforms these features into reconstructed images. The translator is typically a linear regression model trained on paired brain activity and latent feature data. The generator is typically a pre-trained or pre-defined model that generates images from the predicted features. Latent features serve as an intermediate representation that links neural signals and visual content, and recent studies often utilize artificial neural network features (e.g., CNN, CLIP).
In this study, we focus on the translator module, which maps brain activity $\bm{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ to latent features $\bm{y}\in\mathbb{R}^{d_{\mathrm{out}}}$. The training set consists of $n$ paired samples of brain activity and latent features, denoted by $\mathcal{D}=\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$. The latent features $\bm{y}$ are usually high-dimensional representations because they represent a broad spectrum of visual stimuli. Meanwhile, the number of available brain activity samples $n$ is limited; $n$ is much smaller than $d_{\mathrm{out}}$. We therefore define the data scale as the ratio ${n}/{d_{\mathrm{out}}}$, which is typically much smaller than 1 in reconstruction studies. To achieve zero-shot reconstruction, the translator must predict valid latent features for unseen stimuli, even under small data scales.
A common approach to implementing a translator is naive multivariate linear regression, which utilizes a single, shared set of input variables to predict all output dimensions. Representative methods include ridge regression, kernel linear regression, and partial least squares (PLS). These methods differ in how they regularize or project the data, but still rely on shared inputs for all output dimensions.
In contrast, sparse multivariate linear regression allows each output value to select its subset of input variables. According to the classic taxonomy (Guyon and Elisseeff, 2003), selection strategies fall into three groups: wrapper methods (Kohavi and John, 1997), which iteratively search feature subsets by maximizing predictive performance; embedded methods (e.g., Lasso (Tibshirani, 1996), Automatic Relevance Determination (ARD) (MacKay, 1992)), which embed sparsity directly in the training objective; and filter methods (e.g., CFS (Hall, 2000), SIS (Fan and Lv, 2008) ), which rank inputs independently of the learner using simple scores (e.g., correlation, mutual information). In high-dimensional settings such as brain-to-image reconstruction, computational speed is essential, and in practice, the runtime cost scales as wrapper $\gg$ embedded $\gg$ filter.
Previous theoretical work demonstrates that sparsity is particularly beneficial when the sample size is smaller than the dimensionality of the data. For a truly $s$-sparse input-output mapping, information-theoretic bounds show that an ideal estimator can attain near-oracle prediction error once $n\gtrsim s\log(d_{\mathrm{in}}/s)$ (Wainwright, 2009a). The Lasso is the most thoroughly analyzed among practical sparse methods, and theory confirms it reaches essentially the same rate and achieves accurate prediction with $n=\mathcal{O}(s\log(d_{\mathrm{in}}))$ (Donoho, 2006; Wainwright, 2009b). The SIS filter method is also proven to recover the correct support of the mapping with high probability when $d_{\mathrm{in}}$ is exponentially large in $n$ (Fan and Lv, 2008).
3 Output dimension collapse by small data scale
Although both sparse and naive (non-sparse) models are used as the translator module, recent works increasingly rely on the naive model for simplicity and computational efficiency (Seeliger et al., 2018; Ozcelik and VanRullen, 2023; Takagi and Nishimoto, 2023). This section focuses on the naive translator and shows a critical limitation at small data scales.
In brain-to-image reconstruction, the translator must map brain activity $\bm{x}$ onto a high-dimensional image feature space $\bm{y}$ while relying on a limited training sample. If we use a naive multivariate linear regression under such scenarios, a phenomenon “output dimension collapse” (ODC) can occur, whereby the predictions become restricted to the subspace determined by the training outputs (Shirakawa et al., 2025). This issue is especially critical for zero-shot prediction when the training subspace is low-dimensional, because the model inevitably compresses unfamiliar features into that subspace.
To understand why this collapse occurs, let us examine the predicted features of ridge regression. First, we learn a weight matrix $\hat{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ by minimizing the sum of squared errors plus a regularization term:
$\hat{W}$ $=\underset{W}{\operatorname{argmin}}\sum_{i=1}^{n}\lVert\bm{y}^{(i)}-W^{\top}\bm{x}^{(i)}\rVert^{2}+\lambda\lVert W\rVert^{2}_{F}$ (1)
$=\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}{X_{\mathrm{tr}}}^{\top}{Y_{\mathrm{tr}}}.$ (2)
Here, $X_{\mathrm{tr}}=[\bm{x}^{(1)},\bm{x}^{(2)},\ldots,\bm{x}^{(n)}]^{\top}\in\mathbb{R}^{n\times d_{\mathrm{in}}}$ and $Y_{\mathrm{tr}}=[\bm{y}^{(1)},\bm{y}^{(2)},\ldots,\bm{y}^{(n)}]^{\top}\in\mathbb{R}^{n\times d_{\mathrm{out}}}$ are matrices which arrange the brain activity data and latent features of the training data row-wise, respectively. For a test input brain activity $\bm{x}_{\mathrm{te}}$, the prediction is
$\hat{\bm{y}}_{\mathrm{te}}$ $={\hat{W}}^{\top}\bm{x}_{\mathrm{te}}$ (3)
$={Y_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}\bm{x}_{\mathrm{te}}$ (4)
$=Y_{\mathrm{tr}}^{\top}\bm{m}=\sum_{i=1}^{n}m_{i}\bm{y}^{(i)}$ (5)
where $\bm{m}={X_{\mathrm{tr}}}\left({X_{\mathrm{tr}}}^{\top}{X_{\mathrm{tr}}}+\lambda{I}\right)^{-1}\bm{x}_{\mathrm{te}}\in\mathbb{R}^{n}$, and $m_{i}\in\mathbb{R}$ is the $i$-th element of $\bm{m}$. Clearly, the prediction $\hat{\bm{y}}_{\mathrm{te}}$ lies in the subspace spanned by the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. In other words, no matter what the test input is, the predicted output is constrained to that subspace (Figure 2A). Importantly, this property is not the familiar one that confines predictions to the span of the training inputs $\left\{\bm{x}^{(i)}\right\}_{i=1}^{n}$, but instead that it confines predictions to the span of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. This subspace is determined solely by the training outputs and does not depend on the input brain data; given the same training outputs, the same property holds regardless of measurement modality (e.g., fMRI, MEG), subject, or decoding task (e.g., perception, imagery, or illusion). Moreover, this property is not limited to ridge regression; it also holds for other naive multivariate linear regression methods, such as kernel linear regression and PLS (see A.1 for details).
Figure 2: Output dimension collapse. (A) The predictions become restricted to a low-dimensional subspace determined by the training outputs. (B) The best prediction within the training feature subspace provides the lower bound for prediction error in naive multivariate linear regression.
A key condition that leads to output dimension collapse is when the training outputs lie in a low-dimensional subspace. Shirakawa et al. (2025) showed that such a collapse occurs when the training data lack sufficient diversity and the latent features exhibit a cluster structure. In such scenarios, the predicted outputs are pulled toward the clusters in the training data, making generalization to unseen stimuli difficult.
In contrast, we focus here on a different but equally critical factor: small data scales. Even in the absence of explicit cluster structure, having $n<d_{\mathrm{out}}$ inherently forces all predictions into an at-most $n$-dimensional subspace. As $n/d_{\mathrm{out}}$ grows smaller, this low-dimensional subspace becomes more limiting, producing severe output dimension collapse. Consequently, regardless of whether the training outputs form visible clusters, small data scales alone can lead to the same fundamental difficulty in zero-shot prediction.
From another perspective, output dimension collapse due to small data scale is an inevitable consequence of reducing the number of free parameters in the model. The model must learn a weight matrix with many parameters to predict high-dimensional latent features, but the available training samples are insufficient to identify all of them. Therefore, the model must reduce the effective number of free parameters by imposing additional constraints. Naive linear regression accomplishes this by confining all predictions to the linear span of the training outputs. This implicit constraint is precisely what leads to output dimension collapse.
To formally characterize this collapse, we derive the best prediction attainable within the collapsed subspace. This best prediction establishes a lower bound for prediction error under output dimension collapse. Let $\mathcal{A}=\left\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\right\}$ denote the subspace defined by linear combinations of the training outputs $\left\{\bm{y}^{(i)}\right\}_{i=1}^{n}$. The vector $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\in\mathcal{A}$ that is closest to the true latent feature vector $\bm{y}_{\mathrm{te}}$ can be derived analytically as:
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ (6)
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}$ (7)
where $A^{+}$ denotes the Moore-Penrose pseudo-inverse of $A$ (see A.2 for details). This expression shows that $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ is the orthogonal projection of the true latent feature vector onto the subspace $\mathcal{A}$ (Figure 2B).
Crucially, the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ provides the lower bound for prediction error in naive multivariate linear regression:
$\forall\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A},\quad\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}\leq\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}.$ (8)
This inequality shows that no prediction $\hat{\bm{y}}_{\mathrm{te}}$ can ever outperform the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$. Moreover, the prediction error $\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ can be decomposed into two components:
$\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}=\lVert\hat{\bm{y}}_{\mathrm{te}}-\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\rVert^{2}+\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}.$ (9)
The first term $\lVert\hat{\bm{y}}_{\mathrm{te}}-\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}\rVert^{2}$ corresponds to the within-subspace error, and the second term $\lVert\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\rVert^{2}$ corresponds to the out-of-subspace error. When $n<d_{\mathrm{out}}$, the out-of-subspace error is generally non-zero, implying an irreducible error dictated by the training outputs. This error originates from the training outputs being insufficient to cover the universe of possible images. We later verify empirically that this irreducible error is large enough to degrade the quality of reconstructed images, and actual predictions from brain activity are close to this analytic limit.
4 Empirical study: ODC on real data
In the previous section, we showed that naive multivariate linear regression traps in output dimension collapse at small data scales. This section examines how this collapse affects brain-to-image reconstruction in practice. Specifically, we first analyze the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ under various data scales and then compare $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the actual prediction from brain $\hat{\bm{y}}_{\mathrm{te}}$.
Our experiments mainly follow the method of Shen et al. (2019), which employs a Translator–Generator framework to reconstruct images from brain activity. The original method adopts a sparse linear regression model as the translator, and the generator iteratively optimizes the reconstructed image via a CNN-based process. Here, we replace the sparse translator with naive multivariate linear regression to observe the effect of output dimension collapse. For the generator, we adopt Deep Image Prior (DIP) (Ulyanov et al., 2018) as the image prior to visualize the latent features with minimal prior influence, following the protocol of Shen et al. (2019) for iterative image optimization. We also evaluated the method of Ozcelik and VanRullen (2023), which adopts naive multivariate linear regression as the translator (Figure 2). We selected the Deeprecon dataset (Shen et al., 2019), designed explicitly for brain-to-image reconstruction. The training set comprises 1,200 natural images from ImageNet (Deng et al., 2009), and the test set consists of 50 natural images from different categories (unseen during training) and 40 artificial shapes drawn entirely from outside the training domain. Zero-shot prediction beyond the training stimuli is crucial in brain-to-image reconstruction research, so we choose a dataset whose training and test partitions are strictly separated, enabling a rigorous zero-shot evaluation.
To see how the data scale drives output dimension collapse, we progressively reduce the training samples from 1,200 down to 600, 300, and 150, and we compute the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ in each condition. Figure 3A plots the best prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across these different values of $n$. As $n$ decreases, the error grows, indicating that the dimension of the training subspace is shrinking and cannot fully capture the variability in the true latent features. Figure 3B shows example reconstructions generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$ and the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ at each data scale. While the reconstruction from the uncollapsed true feature $\mathcal{G}(\bm{y}_{\mathrm{te}})$ is nearly perfect, the quality of the reconstruction from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ worsens noticeably as $n$ decreases. These results suggest that in reconstruction methods using naive multivariate linear regression, the data scale critically affects reconstruction quality through output dimension collapse.
Figure 3: The best predictions for different training data sizes. (A) The best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across training data sizes 1200, 600, 300, and 150. The vertical axis represents the error of the best prediction, and the horizontal axis represents the sample index sorted by the error at n=n=150. (B) Representative reconstructions from the best predictions for different training data sizes. From top: ground truth, reconstructed images generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best prediction values $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ using 1200, 600, 300, and 150 samples. The left four columns show natural images, and the right two columns show artificial shape images.
To assess how output dimension collapse limits actual prediction performance, we use the full set of 1,200 training images and compare the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the prediction from brain activity $\hat{\bm{y}}_{\mathrm{te}}$. Figure 4A compares the prediction errors of the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the brain prediction $\hat{\bm{y}}_{\mathrm{te}}$. For natural images, the best prediction error accounts for 70% of the total prediction error, and even for artificial images, the proportion reaches 40% (left panel). Examining individual samples reveals a strong correlation between the best and brain prediction errors (right panel). Samples that lie farther from the subspace (i.e., with larger best prediction errors) exhibit correspondingly larger total prediction errors. These observations show that a substantial portion of the prediction error corresponds to the irreducible out-of-subspace error, i.e., a direct consequence of output dimension collapse. Figure 1 presents the corresponding results with subsampled training sets. Figure 4B shows the reconstructed images generated from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ and the brain prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Both reconstructions are similarly blurred, markedly different from the “true” reconstruction. We observe similar results with another reconstruction method (Ozcelik and VanRullen, 2023) (Figure 2). As a reference point, Shen et al.’s original method makes different design choices (e.g., a sparse translator) and can yield sharper images.
Figure 4: Comparison of the best prediction and the brain prediction. (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. The bottom row shows the latest method reconstructions (Shen et al., 2019). The left four columns show natural images, and the right two columns show artificial shape images.
These results confirm that naive multivariate linear regression suffers severely from output dimension collapse due to small data scales. Even the best possible subspace prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ deviates substantially from the true latent $\bm{y}_{\mathrm{te}}$, creating an irreducible error that impedes zero-shot reconstruction. In turn, actual predictions $\hat{\bm{y}}_{\mathrm{te}}$ inherit this limitation and produce blurred reconstructions. These results underscore that naive regression is incompatible with the fundamental structure of reconstruction data: the vast image space and the limited training data.
5 Leveraging sparse brain-to-feature structure
In the previous sections, we demonstrated that naive multivariate linear regression suffers from output dimension collapse at small data scales, restricting predictions to the low-dimensional subspace spanned by the training outputs, thereby limiting generalization to unseen stimuli. This limitation arises from the data structure of brain-to-image reconstruction: the vast image space and the limited training data. In contrast, brain-to-feature mapping is known to be sparse, and several studies have reported improved prediction performance by exploiting this sparsity. In this section, we demonstrate that sparse linear regression, which leverages this structure, can yield low prediction errors even at small data scales. We first show that sparse regression is not subject to output dimension collapse. We then formalize a data generation process and estimate the prediction error characterized by a selection rate $p_{\mathrm{sel}}$ and a hit rate $\pi$.
In brain-to-image reconstruction, employing “brain-like” latent features renders the brain-to-feature mapping sparse. Neurophysiology shows that the visual cortex exhibits local or selective activity in response to visual stimuli (Hubel and Wiesel, 1962; De Valois et al., 1982; Vinje and Gallant, 2000). Correspondingly, several reconstruction studies have adopted latent representations that exhibit similar locality and selectivity, such as local bases or intermediate CNN features (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019). Under these “brain-like” features, only a small subset of voxels should influence each latent dimension, so the input-output mapping becomes sparse (Figure 5A).
Figure 5: Sparse brain-to-feature mapping. (A) If the brain and the latent features exhibit local or selective activity, their relationship should become sparse mapping. (B) Data generation process that reflects the sparse structure of the brain-to-feature mapping. The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
Sparse regression models can effectively exploit the sparse brain-to-feature mapping. The model selects a subset of input variables to assign non-zero weights, thereby reducing the effective number of free parameters. Sparse regression encompasses approaches that first perform variable selection and then fit a model (wrapper methods (Kohavi and John, 1997) and filter methods such as CFS (Hall, 2000) and SIS (Fan and Lv, 2008)), as well as approaches that perform model fitting and variable selection jointly (embedded methods such as Lasso (Tibshirani, 1996) and ARD (MacKay, 1992)). Importantly, these models are not subject to output dimension collapse at small data scales ($n<d_{\mathrm{out}}$). Naive regression uses a single shared set of input variables to predict all output dimensions, which implies $\mathrm{rank}(\hat{W})\leq n$. In contrast, a sparse model can select a different subset of input variables for each output dimension, allowing $\mathrm{rank}(\hat{W})>n$. Conceptually, ODC results from a mismatch between the implicit parameter reduction of the naive model and the data structure. When the true mapping is sparse, explicit reduction through variable selection that matches this structure should be preferred. Theoretically, sparse regression can achieve accurate prediction with small data scales by leveraging sparsity (Wainwright, 2009a; Donoho, 2006; Fan and Lv, 2008). Empirically, reconstruction methods that employ brain-like latent features and sparse regression models have achieved accurate zero-shot reconstruction even with limited training data (Miyawaki et al., 2008; Shen et al., 2019).
To engage in a deeper discussion, we formalize a data generation process that reflects the sparse structure of the brain-to-feature mapping (Figure 5B). Let us assume that the brain activity $\bm{x}\in\mathbb{R}^{d}$ and latent features $\bm{y}\in\mathbb{R}^{d}$ have the same dimensionality $d$ for simplicity. Each sample pair $\left(\bm{x}^{(i)},\bm{y}^{(i)}\right)$ is generated as follows:
$\bm{y}^{(i)}$ $=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)},$ (10)
where the brain activity $\bm{x}^{(i)}$ is generated by sampling $\bm{x}^{(i)}\sim\mathcal{N}(\bm{0},I_{d})$, the noise $\bm{\epsilon}^{(i)}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{d})$, and the teacher weight matrix $W\in\mathbb{R}^{d\times d}$ is column-wise $s$-sparse:
$W_{j,k}=\begin{cases}\frac{1}{\sqrt{s}}&j\in S_{k},\\ 0&j\notin S_{k},\end{cases}\quad(j=1,2,\ldots,d),$ (11)
with $S_{k}\subseteq\{1,2,\ldots,d\}$ a random subset of size $s$ chosen uniformly at random from the $d$ indices. We denote the non-zero weight ratio by $a=s/d$, reflecting the fraction of input dimensions that contribute to each output dimension. The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
Under this generative setting (Figure 5B), we analyze the prediction error of a generic sparse-regression framework trained on the $n$ training pairs $\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$. The model first selects a subset of the inputs and then fits a ridge estimator with an optimally tuned penalty on that support. Let the selection rate $p_{\mathrm{sel}}\in(0,1]$ be the fraction of the $d$ inputs that are selected, and the hit rate $\pi\in[0,1]$ be the fraction of the $s$ truly relevant inputs that are selected. The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: as the sample size $n$ and the dimensionality $d$ both diverge ($n,d\to\infty$) with the data scale $n/d=\gamma<\infty$, the prediction error $R=\frac{1}{d}\mathbb{E}\bigl[\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}\bigr]$ is given by
$R\approx\sigma_{\mathrm{eff}}^{2}\Bigl\{1+\frac{1}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr]\Bigr\},$ (12)
with
$\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2},\qquad\tau^{2}=\frac{\pi}{\sigma_{\mathrm{eff}}^{2}},\qquad\rho=\frac{p_{\mathrm{sel}}}{\gamma}$ (13)
(see A.3 for the derivation). Throughout the paper, the relation $\approx$ indicates asymptotic equality as $n,d\to\infty$ with $n/d<\infty$. In case $p_{\mathrm{sel}}=1$ and $\pi=1$ (all $d$ inputs selected), this expression reduces to the prediction error of ridge regression without variable selection. This expression allows us to estimate the prediction error $R$ from the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for a broad class of sparse regression methods, including wrapper and filter methods. Fully embedded methods such as Lasso (Tibshirani, 1996) or ARD (MacKay, 1992) do not satisfy this framework exactly, but refitting ridge on the selected support aligns them with this framework and is known to yield similar prediction performance (Belloni and Chernozhukov, 2013). Figure 6 shows heatmaps of the prediction error $R$ for sparse regression across combinations of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$. With $p_{\mathrm{sel}}$ fixed, $R$ decreases as $\pi$ increases. Along the diagonal $p_{\mathrm{sel}}=\pi$ (random selection), the error does not decrease, whereas variable selection that attains a high hit rate with a small $p_{\mathrm{sel}}$ does reduce the error.
Figure 6: Prediction error of sparse regression. Each heatmap shows the prediction error $R$ of sparse regression with data scale $n/d=0.2$ (left) and $n/d=0.8$ (right). The horizontal axis represents selection rate $p_{\mathrm{sel}}$, and the vertical axis represents hit rate $\pi$. The diagonal line indicates the case where $p_{\mathrm{sel}}=\pi$, which corresponds to random selection. The upper-left shaded region defined by $\pi\geq p_{\mathrm{sel}}/a$ represents the $(p_{\mathrm{sel}},\pi)$ values that are, in principle, unattainable when $a=0.08$.
Sparse regression reduces prediction error by leveraging the sparse brain-to-feature mapping, a structure that arises when brain-like latent features are employed. By formalizing this sparse structure, we have shown that appropriate variable selection can lead to accurate prediction even at small data scales.
6 Quantifying hit rate and prediction error in practical sparse regression
In the previous section, we analyzed how the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ jointly determine the prediction error in sparse regression. However, the selection algorithm and the data determine these two quantities. In this section, we focus on a concrete variable selection procedure and derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. Furthermore, by comparing the prediction errors by our theory with those from simulations, we confirm the accuracy of the analysis and its applicability.
Among the various variable selection methods, we focus on the simple correlation-based (akin to SIS (Fan and Lv, 2008)) as a concrete method (Figure 7). This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data. For each input vector $\mathbf{x}_{j}=\left(X_{\mathrm{tr}}\right)_{:,j}\in\mathbb{R}^{n}$ and output vector $\mathbf{y}_{k}=\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ across the $n$ training samples, the model computes the sample Pearson correlation coefficient:
$r_{jk}=\frac{(\mathbf{x}_{j}-\hat{\mu}_{\mathrm{x}_{j}})^{\top}(\mathbf{y}_{k}-\hat{\mu}_{\mathrm{y}_{k}})}{n\,\hat{\sigma}_{\mathrm{x}_{j}}\hat{\sigma}_{\mathrm{y}_{k}}},$
where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations. Once the model has computed the correlations $r_{jk}$ for all input-output pairs, it selects the top $p_{\mathrm{sel}}$ fraction of input variables with the highest absolute correlations $|r_{jk}|$ for each output dimension $k$. A separate ridge regression uses only the selected inputs to predict each output dimension. Applying this procedure to all $k$ yields a column-wise sparse support, which may differ across outputs.
Figure 7: Correlation-based filter. The depth labeled $n$ is the sample axis: each slice represents one training sample. For each feature in $y$, compute its correlation with all inputs in $x$ across the $n$ samples, select the top $d_{\mathrm{sel}}$ by absolute correlation, and fit the model using only the selected inputs.
Under the data generation process introduced in Section 5 (Figure 5B), we can derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. We obtain the following relation for large $d$:
$\pi\approx\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)+\alpha\right)+\Phi\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)-\alpha\right),$ (14)
with
$\pi_{-}=\frac{p_{\mathrm{sel}}-a\pi}{1-a},\qquad\alpha=\sqrt{\frac{\gamma}{a(1+\sigma^{2})}},$ (15)
where $\Phi$ is the cumulative distribution function of the standard normal distribution. A detailed derivation is provided in A.4. This relation defines a fixed-point equation for the hit rate $\pi$; we can estimate $\pi$ by numerically solving (14) for $\pi$ given the selection rate $p_{\mathrm{sel}}$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. Figure 8 visualizes this relation for two data scales, $n/d=0.2$ and $n/d=0.8$. For a given selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$ increases as the non-zero weight ratio $a$ decreases, and a larger data scale $n/d$ leads to a larger hit rate $\pi$. In particular, when $a$ is small (i.e., the mapping is sparse), high hit rates $\pi$ are attainable even at small $p_{\mathrm{sel}}$.
Figure 8: Hit rate versus selection rate with a correlation-based filter. Each curve shows the estimated prediction error $R$ for different non-zero weight ratios $a$. The horizontal axis shows $p_{\mathrm{sel}}$, and the vertical axis shows $\pi$. Left: data scale $n/d=0.2$, right: data scale $n/d=0.8$.
We next theoretically estimate the prediction error of two models: naive linear regression, which does not perform variable selection, and sparse linear regression, which performs variable selection using a correlation-based filter. Under the assumed data generation process (Figure 5B), we vary the non-zero weight ratio $a$ and the data scale $n/d$, and estimate the corresponding prediction error by using (12) and (14). We fix the noise level at $\sigma=0.1$ to facilitate comparison across conditions. For the sparse model, the selection rate $p_{\mathrm{sel}}$ is set to minimize the estimated prediction error $R$ under (12) and (14). Figure 9 summarizes the theoretical estimates. For naive linear regression, the prediction error decreases steadily with the data scale $n/d$ but remains insensitive to the sparsity $a$, so a single curve suffices; this confirms that the naive model cannot exploit the sparse structure of the data. For sparse linear regression, the prediction error decreases as the non-zero weight ratio $a$ becomes smaller. These results demonstrate that a practical sparse regression model can translate the sparse structure of the data into substantial accuracy gains and that, when the mapping is sufficiently sparse, it can achieve accurate zero-shot prediction even at small data scales.
Figure 9: The estimated prediction error from theory. The black line shows the prediction error of naive linear regression and is independent of the non-zero weight ratio $a$, so only one curve is displayed. The colored lines show the prediction error of sparse regression for different non-zero weight ratios $a$. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the estimated prediction error $R$.
To assess the accuracy and applicability of our theory, we conduct simulations under four data generation processes (Figure 10). Specifically, we consider: (A) a baseline that exactly matches the assumptions used in our analysis (Figure 5B); (B) Gaussian-distributed non-zero teacher weights that mimic uneven brain-to-feature mapping; (C) correlated signals generated with a Toeplitz covariance with parameter $\rho=0.5$ to reflect correlation structure in neural activity; and (D) input noise representing neural variability and measurement noise (see A.5 for details). Across the baseline, Gaussian weights, and input-noise settings, the theoretical estimates closely match the simulated prediction errors. For the correlated-signal setting, the theory and simulations align well within a practical range of data scales $n/d$, but we observe an increase in error at large data scales. This error increase is plausibly because a correlated signal makes variable selection more difficult, so the selection rate $p_{\mathrm{sel}}$ that is optimal in the uncorrelated case fails to select the correct variables under correlation. These results indicate that our framework provides reliable and informative guidance in practical settings.
Figure 10: The simulated prediction error under four data-generating processes. Each plot shows the simulated prediction error $R$ of sparse regression under four different data-generating processes: (A) Baseline (the same as the setting in our theory), (B) Gaussian-distributed non-zero weights, (C) Correlated signals modeled by a Toeplitz covariance with $\rho=0.5$, (D) Input noise. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the simulated prediction error $R$. Each colored line shows the prediction error for a different non-zero weight ratio $a$.
7 Discussion
In this study, we theoretically analyzed how translator models behave under small data scales. We first clarified how small data scales induce output dimension collapse in naive multivariate linear regression, and we evaluate its impact on the reconstruction pipeline that employs a naive translator. We then mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters. Our results show that a sparse model can achieve accurate prediction even at small data scales by exploiting the sparse brain-to-feature structure. These findings imply that sparse regression does not suffer from output dimension collapse; its gains do not arise from improving prediction within the training subspace but from enabling predictions that extend beyond that subspace.
In Section 3, we derived the theoretical best prediction for naive multivariate linear regression in the translator. Although this best prediction minimizes latent-feature error, it does not always yield the most accurate reconstruction. Even so, it is evident that the output dimension collapse caused by small data scales prevents the model from accurately capturing the true latent features, thereby compromising the quality of the reconstructed images. Moreover, the best prediction analysis is valuable not only for characterizing ODC on small data scales but also for diagnosing ODC by the dataset bias described by Shirakawa et al. (2025). The calculation depends only on the training and test images (no brain data is required), so it can be performed prior to any neuroimaging experiment. Researchers can confirm and, if necessary, adjust their stimulus sets before collecting costly neural recordings.
In Section 5 and Section 6, we evaluated sparse linear regression under the idealized assumptions that model real-world data. Although real datasets do exhibit a performance boost from sparse regression, the improvement is typically more modest than in our simulations. We performed additional simulations under a range of more realistic conditions, yet some discrepancies with actual data are likely to remain. First, our simulations cover relatively high SNR regimes. Today’s brain data contains more noise (Allen et al., 2022; Hebart et al., 2023), suggesting that the effective noise level is higher than the values used in our simulations. Although our theory is applicable regardless of the noise level, obtaining a small prediction error requires sufficiently low noise. Second, the latent features used in practice may not fully satisfy the sparsity assumption. While our simulations considered several sparsity levels, it is challenging to estimate the degree of sparsity in real data. Such sparsity can be expected only when the latent features are truly “brain-like”; yet the latent features used in practice, especially fully connected layers of CNNs or the CLIP embedding, probably do not entirely meet this ideal. Improvements in neural recording denoising techniques and the design of AI feature spaces that empirically approximate the brain’s representations will likely enable zero-shot reconstruction at small data scales.
Although we examined only linear translators, recent reconstruction methods sometimes deploy nonlinear architectures as the translator module (Chen et al., 2023; Scotti et al., 2023, 2024). For instance, Chen et al. (2023) employed Masked Brain Modeling, embedding structural priors of brain activity into the translator in an attempt to improve prediction accuracy. Although nonlinear networks can encode such inductive biases, output dimension collapse remains a concern. In the neural-tangent-kernel (NTK) regime (Jacot et al., 2018), nonlinear models behave like kernel linear models, so they cannot avoid output dimension collapse. Empirically, Shirakawa et al. (2025) reported that nonlinear translators used for reconstruction also experience collapse-like phenomena. A rigorous theoretical analysis of whether nonlinear translators can genuinely support zero-shot prediction remains important for future work.
Furthermore, while our study focused on the translator, the generator module might partially compensate for inaccuracies in latent feature prediction. In the latest method, as described in Section 3, the improved visual appearance arises not only from replacing naive linear regression with sparse regression but also from generator-side refinements, such as stronger image priors, latent feature scaling, and carefully chosen feature inversion losses. More recently, some reconstruction pipelines began to adopt diffusion models as the generator (Ozcelik and VanRullen, 2023; Cheng et al., 2023; Scotti et al., 2024). These models enable us to generate photorealistic images even from inaccurate latent features. However, this approach carries the risk of introducing information not actually present in the brain’s representation. In other words, if the generator adds unnecessary details, the reconstructed image may stray from the accurate perceptual content encoded in neural activity. The “Recovery Check” ensures that only the information contained in the latent features is reconstructed (Shirakawa et al., 2025).
Finally, our results apply to data-limited prediction problems at the interface of neuroscience and machine learning, including brain decoding and brain encoding across vision, speech, and language. For example, Brain-Score assesses brain-DNN alignment by fitting linear encoding models from model features to neural responses (Schrimpf et al., 2020). When solving regression at small data scales, output dimension collapse is a critical concern. Sparse regression is effective in avoiding collapse; however, in encoding settings, the selected DNN features tend to exhibit substantial overlap across voxels (Nonaka et al., 2021). Although we do not treat this situation here, determining how such selection concentration shapes prediction remains an important direction for future work.
7.1 Conclusion
The significance of our study lies in its dual contribution: First, we show that naive linear regression suffers from output dimension collapse at small data scales, restricting generalization beyond the training data. Second, we demonstrate that sparse linear regression can achieve accurate prediction at small data scales by mathematically characterizing the prediction error. These findings provide a practical guideline for advancing image reconstruction without relying on large-scale data.
8 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT (GPT-4o, o3, 5) in order to improve readability and English translation. After using these tools/services, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
9 Acknowledgements
We thank our laboratory team, especially Ken Shirakawa and Hideki Izumi, for their invaluable feedback and insightful suggestions on the manuscript. This work was supported by Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705, JP25H00450 to Y.K., and JP21K17821 to Y.N.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, JPMJCR22P3 to Y.K., and BOOST grant JPMJBS2407 to K.O.), New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.), and Guardian Robot Project, RIKEN.
Appendix A Supplementary Materials
A.1 Output subspace restriction in kernel linear regression and PLS
This section shows that kernel linear regression and PLS produce predictions that are linear combinations of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$.
Kernel linear regression.
Let $k(\cdot,\cdot)$ be a positive semidefinite kernel and $K\in\mathbb{R}^{n\times n}$ the Gram matrix with entries $K_{ij}=k(\bm{x}^{(i)},\bm{x}^{(j)})$. For a test input $\bm{x}_{\mathrm{te}}$, define $\bm{k}_{\mathrm{te}}=[k(\bm{x}_{\mathrm{te}},\bm{x}^{(1)}),\ldots,k(\bm{x}_{\mathrm{te}},\bm{x}^{(n)})]^{\top}$. Kernel linear regression fits functions in the RKHS and, by the representer theorem, the multi-output predictor takes the matrix form:
$\hat{\bm{y}}_{\mathrm{te}}\;=\;A^{\top}\bm{k}_{\mathrm{te}}\quad\text{with}\quad A\;=\;(K+\lambda I)^{-1}Y_{\mathrm{tr}}.$
Therefore
$\hat{\bm{y}}_{\mathrm{te}}=Y_{\mathrm{tr}}^{\top}(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}=Y_{\mathrm{tr}}^{\top}\bm{m}=\sum_{i=1}^{n}m_{i}\,\bm{y}^{(i)},$
where $\bm{m}=(K+\lambda I)^{-1}\bm{k}_{\mathrm{te}}\in\mathbb{R}^{n}$. Thus, every prediction is a linear combination of the training outputs, independent of the particular choice of kernel.
Partial least squares (PLS).
We consider multivariate PLS with $r$ components on $(X_{\mathrm{tr}},Y_{\mathrm{tr}})$. Two standard variants are NIPALS and SIMPLS, which differ in how they construct the training score matrix $T\in\mathbb{R}^{n\times r}$ and the out-of-sample score map $t(\bm{x}_{\mathrm{te}})\in\mathbb{R}^{r}$, yet the prediction formula below does not depend on which algorithm produced $T$ and $t(\cdot)$.
Given $T$, the Y-side loading follows from least squares:
$Q^{\top}\;=\;\underset{B\in\mathbb{R}^{r\times d_{\mathrm{out}}}}{\operatorname{argmin}}\|\,Y_{\mathrm{tr}}-TB\,\|_{F}^{2}\;=\;(T^{\top}T)^{-1}T^{\top}Y_{\mathrm{tr}}.$
For a test input $\bm{x}_{\mathrm{te}}$, PLS predicts
$\hat{\bm{y}}_{\mathrm{te}}\;=\;t(\bm{x}_{\mathrm{te}})^{\top}Q^{\top}\;=\;t(\bm{x}_{\mathrm{te}})^{\top}(T^{\top}T)^{-1}T^{\top}Y_{\mathrm{tr}}\;=\;Y_{\mathrm{tr}}^{\top}\,\bm{m},$
where $\bm{m}\;=\;T(T^{\top}T)^{-1}t(\bm{x}_{\mathrm{te}})\in\mathbb{R}^{n}$. Thus every prediction $\hat{\bm{y}}_{\mathrm{te}}$ is a linear combination of the training outputs $\{\bm{y}^{(i)}\}_{i=1}^{n}$. This conclusion holds for NIPALS and SIMPLS alike, regardless of the specific construction of $T$ and $t(\cdot)$.
A.2 The best prediction
The best prediction in naive multivariate linear regression is derived analytically as follows:
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$
$=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}},$
where $A^{+}$ is the Moore-Penrose pseudo-inverse of $A$.
Proof.
Since $\mathcal{A}=\left\{Y_{\mathrm{tr}}^{\top}\bm{m}\mid\bm{m}\in\mathbb{R}^{n}\right\}$, we can express the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ as
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=\underset{\hat{\bm{y}}_{\mathrm{te}}\in\mathcal{A}}{\operatorname{argmin}}\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}$
$=\underset{\bm{m}\in\mathbb{R}^{n}}{\operatorname{argmin}}\lVert Y_{\mathrm{tr}}^{\top}\bm{m}-\bm{y}_{\mathrm{te}}\rVert^{2}$
The normal equation obtains the solution:
$Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\bm{m}=Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
Thus, the best prediction is
$\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ $=Y_{\mathrm{tr}}^{\top}\left(Y_{\mathrm{tr}}Y_{\mathrm{tr}}^{\top}\right)^{+}Y_{\mathrm{tr}}\bm{y}_{\mathrm{te}}.$
A.3 Prediction error for sparse linear regression
We quantify the prediction error when the sparse regression retains a fraction $p_{\mathrm{sel}}$ of the inputs and captures a fraction $\pi$ of the truly relevant ones. We focus on the asymptotic regime where both the sample size $n$ and the dimensionality $d$ are taken to infinity, while the data scale $n/d$ stays finite. By applying the risk formula for ridge regression by Dicker (2016) to the post-selection problem, we quantify the prediction error of sparse regression.
Consider the linear model that remains after variable selection for a single coordinate $y_{k}$ of $\bm{y}$. Split every input vector $\bm{x}^{(i)}\in\mathbb{R}^{d}$ into two parts: the selected part $\bm{\xi}^{(i)}_{k}\in\mathbb{R}^{d_{\mathrm{sel}}}$ and the unselected part $\bm{\zeta}^{(i)}_{k}\in\mathbb{R}^{d-d_{\mathrm{sel}}}$, and similarly the weight vector $\bm{w}_{k}=W_{:,k}\in\mathbb{R}^{d}$ into the corresponding $\bm{u}_{k}\in\mathbb{R}^{d_{\mathrm{sel}}}$ and $\bm{v}_{k}\in\mathbb{R}^{d-d_{\mathrm{sel}}}$. With these notations
$y_{k}^{(i)}$ $=\bm{w}_{k}^{\top}\bm{x}^{(i)}+\epsilon_{k}^{(i)}$ (16)
$=\bm{u}_{k}^{\top}\bm{\xi}^{(i)}_{k}+\bm{v}_{k}^{\top}\bm{\zeta}^{(i)}_{k}+\epsilon_{k}^{(i)}$ (17)
$=\bm{u}_{k}^{\top}\bm{\xi}^{(i)}_{k}+\eta_{k}^{(i)}$ (18)
where $\eta_{k}^{(i)}=\bm{v}_{k}^{\top}\bm{\zeta}^{(i)}_{k}+\epsilon_{k}^{(i)}$ is the combined noise term. The selected part of weight vector $\bm{u}_{k}$ contains $\pi s$ non-zero entries, and the unselected part $\bm{v}_{k}$ contains $(1-\pi)s$ non-zero entries, hence $\|\bm{u}_{k}\|^{2}=\pi$ and $\|\bm{v}_{k}\|^{2}=1-\pi$. Further, since $\bm{x}^{(i)}$ is generated from $\mathcal{N}(\bm{0},I_{d})$ and independent of $\bm{\epsilon}^{(i)}$, the combined noise satisfies
${\eta}_{k}^{(i)}\sim\mathcal{N}(0,\sigma_{\mathrm{eff}}^{2}),$ (19)
with $\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2}$, and it is independent of $\bm{\xi}^{(i)}_{k}$. The post-selection problem is therefore a standard linear regression with effective dimension $d_{\mathrm{sel}}$ and effective noise variance $\sigma_{\mathrm{eff}}^{2}$.
The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: in the limit $n,d_{\mathrm{sel}}\to\infty$ with $n/d_{\mathrm{sel}}<\infty$, the risk of the ridge estimator with the optimal regularization parameter is
$\mathbb{E}\|\hat{\bm{u}}_{k}-\bm{u}_{k}\|^{2}\approx\frac{\sigma_{\mathrm{eff}}^{2}}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr],$ (20)
with
$\tau^{2}=\frac{\|\bm{u}_{k}\|^{2}}{\sigma_{\mathrm{eff}}^{2}}=\frac{\pi}{\sigma_{\mathrm{eff}}^{2}},\quad\rho=\frac{d_{\mathrm{sel}}}{n}=\frac{p_{\mathrm{sel}}}{\gamma},\quad\sigma_{\mathrm{eff}}^{2}=1-\pi+\sigma^{2},$ (21)
where $\gamma=n/d$ is the data scale. For the test data $(\bm{x}_{\mathrm{te}},\bm{y}_{\mathrm{te}})$, we can split $\bm{x}_{\mathrm{te}}$ into $\bm{\xi}_{\mathrm{te},k}\in\mathbb{R}^{d_{\mathrm{sel}}}$ and $\bm{\zeta}_{\mathrm{te},k}\in\mathbb{R}^{d-d_{\mathrm{sel}}}$ in the same way as above. The $k$-th output coordinate of the test data $\bm{y}_{\mathrm{te}}$ is similarly expressed as
$y_{\mathrm{te},k}=\bm{u}_{k}^{\top}\bm{\xi}_{\mathrm{te},k}+\eta_{\mathrm{te},k},$ (22)
where $\eta_{\mathrm{te},k}\sim\mathcal{N}(0,\sigma_{\mathrm{eff}}^{2})$ is independent of $\bm{\xi}_{\mathrm{te},k}$. The learned weight on the unselected part $\bm{\zeta}_{\mathrm{te},k}$ is zero, so the prediction for the $k$-th output coordinate is
$\hat{y}_{\mathrm{te},k}$ $=\hat{\bm{w}}_{k}^{\top}\bm{x}_{\mathrm{te}}$ (23)
$=\hat{\bm{u}}_{k}^{\top}\bm{\xi}_{\mathrm{te},k},$ (24)
where $\hat{\bm{u}}_{k}$ is the learned weight on the selected part $\bm{\xi}_{\mathrm{te},k}$. Therefore, the prediction error $R$ is
$R$ $=\frac{1}{d}\mathbb{E}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$ (25)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|\hat{y}_{\mathrm{te},k}-y_{\mathrm{te},k}\|^{2}\right]$ (26)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|(\hat{\bm{u}}_{k}-\bm{u}_{k})^{\top}\bm{\xi}_{\mathrm{te}}-\eta_{\mathrm{te},k}\|^{2}\right]$ (27)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|(\hat{\bm{u}}_{k}-\bm{u}_{k})^{\top}\bm{\xi}_{\mathrm{te}}\|^{2}+\|\eta_{\mathrm{te},k}\|^{2}\right]$ (28)
$=\frac{1}{d}\sum_{k=1}^{d}\mathbb{E}\left[\|\hat{\bm{u}}_{k}-\bm{u}_{k}\|^{2}+\sigma_{\mathrm{eff}}^{2}\right]$ (29)
$\approx\sigma_{\mathrm{eff}}^{2}\Bigl\{1+\frac{1}{2\rho}\Bigl[\tau^{2}(\rho-1)-\rho+\sqrt{\bigl(\tau^{2}(\rho-1)-\rho\bigr)^{2}+4\rho^{2}\tau^{2}}\Bigr]\Bigr\}.$ (30)
This expression gives the prediction error of sparse regression with a selection rate $p_{\mathrm{sel}}$ and a hit rate $\pi$.
A.4 Selection rate-hit rate relation for correlation-based filter
We relate the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ of the correlation-based filter to characterize the performance of the practical filter. We work in the same asymptotic regime as in the preceding analysis: $n,d\to\infty$ with $n/d<\infty$. All $\approx$’ relations are asymptotic equalities in this limit.
In the following, we express the column slices of data matrices $\left(X_{\mathrm{tr}}\right)_{:,j},\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ as $\mathbf{x}_{j}$ or $\mathbf{y}_{k}$. We intentionally use different symbol from $\bm{x}^{(i)}$ or $\bm{y}^{(i)}$ which indicates $i$-th single instances. For each output coordinate $k\in\{1,\ldots,d\}$, we can rewrite the data generation process as
$\mathbf{y}_{k}$ $=X_{\mathrm{tr}}W_{:,k}+\mathbf{\epsilon}_{k}$ (31)
$=\sum_{j=1}^{d}W_{jk}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (32)
$=\frac{1}{\sqrt{s}}\sum_{j\in\mathcal{S}_{k}}\mathbf{x}_{j}+\mathbf{\epsilon}_{k},$ (33)
where $W_{:,k}$ is the $k$-th column of the teacher weight matrix $W$, $\mathbf{\epsilon}_{k}\in\mathbb{R}$ is the $k$-th output noise vector, and $\mathcal{S}_{k}=\left\{j\mid W_{jk}\neq 0\right\}$ is the informative index set for the $k$-th output coordinate. Within this data generation process, $\mathbf{x}_{j}\sim\mathcal{N}(\bm{0},I_{n})$ and $\mathbf{\epsilon}_{k}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{n})$ hold.
The correlation-based filter ranks each input variable $j$ according to its absolute correlation with the output $k$:
$r_{jk}=\frac{(\mathbf{x}_{j}-\hat{\mu}_{\mathrm{x}_{j}})^{\top}(\mathbf{y}_{k}-\hat{\mu}_{\mathrm{y}_{k}})}{n\,\hat{\sigma}_{\mathrm{x}_{j}}\hat{\sigma}_{\mathrm{y}_{k}}},$ (34)
where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations. Because $\textbf{x}_{j}\sim\mathcal{N}(\bm{0},I_{n})$ and $\textbf{y}_{k}\sim\mathcal{N}(\bm{0},(1+\sigma^{2})I_{n})$ hold, the means and standard deviations satisfy $\hat{\mu}_{\mathrm{x}_{j}}\approx 0,\hat{\sigma}_{\mathrm{x}_{j}}\approx 1,\hat{\mu}_{\mathrm{y}_{k}}\approx 0,\hat{\sigma}_{\mathrm{y}_{k}}\approx\sqrt{1+\sigma^{2}}$. Thus, the correlation coefficient simplifies to
$r_{jk}\approx\frac{\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle}{n\sqrt{1+\sigma^{2}}}=\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle\times\mathrm{const}.$ (35)
Hence, the ranking by the absolute correlation coefficient $|r_{jk}|$ is almost equivalent to the ranking by the absolute inner product $|z_{jk}|=\left|\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle\right|$.
We first derive the expectation and variance of the inner product $z_{jk}=\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle$ for informative inputs ($j\in\mathcal{S}_{k}$) and uninformative inputs ($j\notin\mathcal{S}_{k}$). For an informative input $j\in\mathcal{S}_{k}$, we can write
$z_{jk}$ $=\langle\mathbf{x}_{j},\mathbf{y}_{k}\rangle$ (36)
$=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (37)
$=\frac{1}{\sqrt{s}}\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (38)
Independence of the Gaussian components yields
$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (39)
$=\frac{n}{\sqrt{s}}+0+0=\frac{n}{\sqrt{s}},$ (40)
$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{j}\rangle]+\frac{1}{s}\sum_{t\in\mathcal{S}_{k}\setminus j}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (41)
$=\frac{2n}{s}+\frac{(s-1)n}{s}+n\sigma^{2}=n(\frac{s+1}{s}+\sigma^{2}).$ (42)
Similarly, for an uninformative input $j\notin\mathcal{S}_{k}$, we find
$z_{jk}$ $=\langle\mathbf{x}_{j},\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbf{x}_{t}+\mathbf{\epsilon}_{k}\rangle$ (43)
$=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle+\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle.$ (44)
$\mathbb{E}[z_{jk}]$ $=\frac{1}{\sqrt{s}}\sum_{t\in\mathcal{S}_{k}}\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathbb{E}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (45)
$=0+0=0,$ (46)
$\mathrm{Var}[z_{jk}]$ $=\frac{1}{s}\sum_{t\in\mathcal{S}_{k}}\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{x}_{t}\rangle]+\mathrm{Var}[\langle\mathbf{x}_{j},\mathbf{\epsilon}_{k}\rangle]$ (47)
$=n(1+\sigma^{2}).$ (48)
Next, we derive the relation between the hit rate $\pi$ and the selection rate $p_{\mathrm{sel}}$ of the correlation-based filter. Fix an output $k$ and let the true support be $\mathcal{S}_{k}=\{j\mid w_{jk}\neq 0\}$ with $|\mathcal{S}_{k}|=s=ad$. For a common threshold $t_{k}>0$, define the selection set $\hat{\mathcal{S}}_{k}(t_{k})=\{j\mid|z_{jk}|>t_{k}\}$. Recall the selection rate $p_{\mathrm{sel}}=\frac{|\hat{\mathcal{S}}_{k}(t_{k})|}{d}=\Pr(j\in\hat{\mathcal{S}}_{k}(t_{k}))$ and the hit rate $\pi=\Pr(j\in\hat{\mathcal{S}}_{k}(t)\mid j\in\mathcal{S}_{k})$. Define the null retention probability $\pi_{-}:=\Pr(j\in\hat{\mathcal{S}}_{k}(t_{k})\mid j\notin\mathcal{S}_{k})$. Noting that $a=\Pr(j\in\mathcal{S}_{k})$, we obtain the identity
$p_{\mathrm{sel}}$ $=a\,\pi+(1-a)\,\pi_{-}.$ (49)
We proceed by approximating $\pi$ and $\pi_{-}$ and then eliminating the threshold. Denote by $\Phi(\cdot)$ and $\Phi^{-1}(\cdot)$ the CDF and inverse-CDF of the standard normal distribution. Writing $u_{k}:=t_{k}/\sqrt{n}$ and $\Delta:=\sqrt{n/s}=\sqrt{\gamma/a}$ with data scale $\gamma=n/d$, we obtain
$\pi$ $=\Pr(|z_{jk}|>t_{k}\mid j\in\mathcal{S}_{k})$ (50)
$=\Phi\!\left(-\frac{t_{k}-\frac{n}{\sqrt{s}}}{\sqrt{\,n\!\left(\frac{s+1}{s}+\sigma^{2}\right)}}\right)+\Phi\!\left(-\frac{t_{k}+\frac{n}{\sqrt{s}}}{\sqrt{\,n\!\left(\frac{s+1}{s}+\sigma^{2}\right)}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right)$ (51)
$=\Phi\!\left(-\frac{u_{k}-\Delta}{\sqrt{\,1+\sigma^{2}+\frac{1}{s}\,}}\right)+\Phi\!\left(-\frac{u_{k}+\Delta}{\sqrt{\,1+\sigma^{2}+\frac{1}{s}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right)$ (52)
$=\Phi\!\left(-\frac{u_{k}-\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\Phi\!\left(-\frac{u_{k}+\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right).$ (53)
(51) follows from Berry-Esseen theorem, and (53) follows from
$\Phi\!\left(-\frac{u_{k}\pm\Delta}{\sqrt{\,1+\sigma^{2}+\frac{1}{s}\,}}\right)=\Phi\!\left(-\frac{u_{k}\pm\Delta}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(1/s\right).$ (54)
Similarly,
$\pi_{-}$ $=\Pr(|z_{jk}|>t_{k}\mid j\notin\mathcal{S}_{k})$ (55)
$=2\,\Phi\!\left(-\frac{t_{k}}{\sqrt{\,n(1+\sigma^{2})\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right)$ (56)
$=2\,\Phi\!\left(-\frac{u_{k}}{\sqrt{\,1+\sigma^{2}\,}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right).$ (57)
Combining (53) and (57) with (49) yields the two-equation relation
$p_{\mathrm{sel}}$ $=a\!\left[\Phi\!\left(-\tfrac{u_{k}-\Delta}{\sqrt{1+\sigma^{2}}}\right)+\Phi\!\left(-\tfrac{u_{k}+\Delta}{\sqrt{1+\sigma^{2}}}\right)\right]$
$\quad+(1-a)\,2\,\Phi\!\left(-\tfrac{u_{k}}{\sqrt{1+\sigma^{2}}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right),$ (58)
$\pi$ $=\Phi\!\left(-\tfrac{u_{k}-\Delta}{\sqrt{1+\sigma^{2}}}\right)+\Phi\!\left(-\tfrac{u_{k}+\Delta}{\sqrt{1+\sigma^{2}}}\right)+\mathcal{O}\!\left(n^{-\frac{1}{2}}\right),$ (59)
which links $p_{\mathrm{sel}}$ and $\pi$ through the common threshold $u_{k}>0$. Eliminating $u_{k}$ gives, for $a\in(0,1)$,
$\pi\approx\Phi\!\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)+\alpha\right)+\Phi\!\left(\Phi^{-1}\left(\frac{\pi_{-}}{2}\right)-\alpha\right)$ (60)
where
$\pi_{-}=\frac{p_{\mathrm{sel}}-a\pi}{1-a},\quad\alpha=\frac{\Delta}{\sqrt{1+\sigma^{2}}}=\sqrt{\frac{\gamma}{a(1+\sigma^{2})}},$ (61)
while in the boundary case $a=1$ the relation reduces to $p_{\mathrm{sel}}=\pi$.
A.5 Detailed Simulation Settings
We conducted simulations under the following four generative conditions:
  • (A)
    Baseline: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{x}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. Each non-zero entry in the teacher matrix $W$ is set to $1/\sqrt{s}$, with the remaining entries set to 0. The input vector $\bm{x}[{(i)}]$ is drawn from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
  • (B)
    Gaussian non-zero weights: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as in (A), but the non-zero entries in the teacher matrix are sampled from a standard normal distribution $\mathcal{N}(0,\frac{1}{s})$. Inputs $\bm{x}^{(i)}$ are independently drawn from $\mathcal{N}(0,1)$, as in (A).
  • (C)
    Correlated signal: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as in (A), with the same binary non-zero structure in the teacher matrix $W$. The input vector $\bm{x}^{(i)}$ is sampled from a zero-mean multivariate normal distribution with a Toeplitz covariance matrix $\Sigma_{ij}=\rho^{|i-j|}$, where $\rho=0.5$.
  • (D)
    Input noise: Each sample pair $(\bm{x}^{(i)},\bm{y}^{(i)})$ is generated as $\bm{y}^{(i)}=W^{\top}\bm{z}^{(i)},\bm{x}^{(i)}=\bm{z}^{(i)}+\bm{\epsilon}^{(i)}$, where $\bm{\epsilon}^{(i)}\sim\mathcal{N}(0,\sigma^{2}I)$. The teacher matrix $W$ has the same binary non-zero structure as in (A). The input vector $\bm{z}^{(i)}$ is sampled from a standard normal distribution with independent components $\mathcal{N}(0,1)$.
The simulations are performed with the following parameters: the dimensionality $d=1000$ and the noise level $\sigma=0.1$.
Appendix B Supplementary Figures
Figure 1: Comparison of the best prediction and the brain prediction for subsampled training sets. Each panel shows the same analysis as in Figure 4 with subsampled training sets of 1200, 600, 300, and 150 samples. Each dot represents one image, and the dotted line indicates where the two errors are equal. Blue for natural images, orange for artificial shapes. As the data scale decreases, the output dimension collapse intensifies, and the brain’s prediction approaches the best prediction.
Figure 2: Comparison of the best prediction and the brain prediction on another reconstruction method (Ozcelik and VanRullen, 2023). (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Similar results are observed: the brain prediction and best prediction are close to each other but far from the true feature.